word problem
Lila, Alice, and Mary are hunting for seashells. Together they find a total of 31 seashells. Alice finds two more than Lila. and Mary finds one less than lila. how many did each girl find?
x=number of seashells
total=31
Alice=n+2 ??
2x=number of shells Lila has
2x-1= Mary??
not sure how to find the variables. your help is much appreciated thanks.

Answer by mananth(16946)

Lila, Alice, and Mary are hunting for seashells. Together they find a total of 31 seashells. Alice finds two more than Lila. and Mary finds one less than lila. how many did each girl find?
L + A +M 31
A=L+2
M=L-1
plug A & M
L+ (L+2)+(L-1)=31
3L+1=31
3L=30
L=10
A=L+2=12